Let f(X) = −log|X| with the domain of f is the set of positive-definite matrices, denoted by S++. Notice that |X| is the determinant of matrix X. Show that f is convex.
Answer #1

since f(x) = -log X

by the definition of convexivity we have to show that

f(xX + (1 - XY) < f(x) + (1 - 1)f(Y)) where 0< \lambda < 1

Now, log XX + (1 - XY > log(4X + (1 - XY) { because, det (A+B) \geq Det(A) +Det (B) }

=log|\lambda X|+log (|(1-\lambda )Y|/|\lambda X|) {USING , log mp = log m +log p}

=(nlog\lambda +log| X|)+(nlog(1-\lambda)/\lambda ) +log (| Y|/| X|)) WHERE n is the order of the matrix

  \geq (\lambda log| X|)+(((1-\lambda )/\lambda )log (| Y|/| X|))

  \geq (\lambda log| X|)+(1-\lambda )log (| Y|/| X|) {because  (1/2-1) > (1 - ) }

> Vlog X + (1 - 1)log Y

therefore -logAX + (1 - 1) <-Nog X - (1 - 1)log|Y|

i.e . f(XX + (1 - XY)<Af(x) + (1 - 1)f(Y)

hence f is convex